OpenStudy (anonymous):

In a parallel circuit, what affect will adding more resistors to the circuit have?

OpenStudy (agent0smith):

More resistors in parallel lowers the overall resistance. You're providing more "pathways" for current to travel through.

OpenStudy (anonymous):

so in other words The current will decrease by a factor proportional to the number of resistors?

OpenStudy (agent0smith):

Current INCREASES. Resistance decreases. V = IR This is the formula you'd use for total resistance. \[\Large \frac{ 1 }{ R_T } =\frac{ 1 }{R_1 }+\frac{ 1 }{R_2 }+\frac{ 1 }{R_3 }+...\frac{ 1 }{R_n }\]

OpenStudy (anonymous):

ohhhhh!!! now I get it ...thank you very much :3

OpenStudy (agent0smith):

@Mateaus the voltage drop is the same across all the resistors.

OpenStudy (agent0smith):

The voltage is the same across every resistor. If you have a 12V battery with five resistors in parallel, each resistor has 12V over it.

OpenStudy (anonymous):

I did understood so that's why ,thanks to both you ^.^)